# SEAL: Towards Safe Autonomous Driving via Skill-Enabled Adversary Learning for Closed-Loop Scenario Generation

### v1

@@ Abstract
Verification and validation of autonomous driving (AD) systems and components is of increasing importance, as such technology increases in real-world prevalence. Safety-critical scenario generation is a key approach to robustify AD policies through closed-loop training. However, existing approaches for scenario generation rely on simplistic objectives, resulting in overly-aggressive or non-reactive adversarial behaviors. To generate diverse adversarial yet realistic scenarios, we propose SEAL, a scenario perturbation approach which leverages learned scoring functions and adversarial, human-like skills. SEAL-perturbed scenarios are more realistic than SOTA baselines, leading to improved ego task success across real-world, in-distribution, and out-of-distribution scenarios, of more than 20%. To facilitate future research, we release our code and tools: https://github.com/cmubig/SEAL
@@ I Introduction
Scenarios are often curated in the form of large datasets of real-world recorded driving traces, providing a basis for assessing human behaviors and for training machine learning models [8, 9, 10]. AD subsystems are then asked to perform tasks such as forecasting the future motion of various road users or controlling the behavior of a particular vehicle in a simulated reconstruction [11, 12, 13, 14]. However, the presence of critical scenarios in collected datasets is exceedingly low, a problem identified as the “curse-of-rarity” in autonomous driving [15, 16, 17]. Thus, programmatically generating safety-critical scenarios has become necessary. To ensure that generated scenarios retain realistic properties, it is appealing to perturb the behavior of one or more agents in a principled way, rather than using first principles to painstakingly assemble a scenario from scratch [18, 19, 20, 21]. In this setting, one agent is referred to as the ego agent, while the modified background traffic participants are adversary agent(s), who attempt to attack the ego in some way.
@@
Therefore, in this paper, we propose and evaluate a method for Skill-Enabled Adversary Learning (SEAL), which yields significantly improved downstream ego behavior, in closed-loop training with safety-critical scenario generation. Our method addresses the identified limitations in prior art by introducing two novel components, as shown in LABEL:fig:overview. First, we introduce a learned scoring function to anticipate how a reactive ego agent will respond to a candidate adversarial agent behavior. We quantify both collision closeness and induced ego behavior deviation, thus providing a broadened understanding of safety criticality. Second, we develop a reactive adversary policy; in particular, inspired by human cognition, we leverage a hierarchical framework that is akin to how humans operate vehicles [26] and we create an adversarial prior that selects human-like skill primitives to increase criticality while maintaining realism.
Furthermore, we argue that safety-critical scenario generation approaches should be evaluated in terms of behavior realism and usability in improving ego policy development, rather than induced criticality alone. Much prior work, however, relies on evaluating ego policies by leveraging their scenario generation approach on a set of base scenarios, wherein safety-critical behavior is effectively in-distribution with respect to the training examples and the family of heuristic scene perturbations [21, 27]. Thus, we extend recent work on scenario characterization [16] to instead identify real (non-generated) yet still safety-relevant scenarios in datasets, thereby creating a realistic, out-of-distribution evaluation setting. We argue that while ego performance on in-distribution generated scenes is informative, performance on real challenging scenes ultimately matters the most.
@@
    We propose two novel techniques for safety-critical perturbation: (i) a learned scoring function, to select over candidate trajectories; and (ii) a reactive, adversarial skill policy for increased realism in adversary behavior.
@@ II-A Scenario Generation in Autonomous Driving
Approaches for generating scenarios that reproduce the distribution of normal driving behavior have been extensively explored. Some methods ensure the diversity of generated traffic behavior [28, 29], while others aim for controllability through rule-based or language-driven specifications [30, 31, 32]. However, due to the rarity of safety-critical events in recorded data [15, 16, 17], other approaches have focused on directly generating corner-case scenarios by injecting adversarial behaviors. Earlier works in safety-critical scenario generation relied on gradient-based optimization approaches that require access to vehicle dynamics [27, 24, 33], a limitation in model-free settings. Other methods, such as diffusion-based approaches [25, 34], are compute-intensive and impractical to be used in a closed-loop manner. Efficient methods like CAT [21] and GOOSE [35], which leverage trajectory prediction priors and reinforcement learning (RL) respectively, are non-reactive to the ego agent and often prioritize simple collision objectives. In contrast, our approach efficiently generates reactive, nuanced adversarial behavior, providing a stronger training signal and improving closed-loop agent performance.
@@ II-B Robust Training and Evaluation in Autonomous Driving
Several techniques for robustifying AD policies against safety-critical and out-of-distribution scenarios have been explored. Formal methods, such as Hamilton-Jacobi (HJ) reachability, have been utilized in various driving tasks, but struggle with dimension scaling [36, 37]. Similarly, domain randomization has been used as a form of data augmentation (e.g., randomizing vehicle control parameters [38] or scenario initial states [39]) but requires excessive sampling to cover a sufficient domain size. Thus, adversarial training has been increasingly used, either as a fine-tuning scheme [25, 24] or in a fully closed-loop training pipeline [21, 33], providing adaptive, continuous feedback to an ego agent.
@@ IV Approach: Skill-Enabled Adversary Learning for Scenario Generation
To increase the criticality while maintaining the realism of a scenario, we propose the Skill-Enabled Adversary Learning (SEAL) approach for perturbation-based scenario generation. We share a similar insight to CAT [21] of utilizing a probabilistic trajectory prediction model, $\pi_{\text{gen}}$, to produce diverse candidate future paths that an adversary could take, conditioned on $\mathbf{S}$. However, there are several challenges with directly providing a future path for the adversary to follow from $\pi_{\text{gen}}(\mathbf{S})$:
@@
To address these challenges, we introduce two key components: a learned scoring function to select trajectories more flexibly than heuristic methods, and an adversarial skill policy that enables more reactive and human-like behavior.
@@
IV-A Learned Scoring Function
@@ IV-B Adversarial Skill Learning
We design a reactive policy, $\pi_{\text{adv}}$, to guide the adversary’s behavior, unlike recent works [21, 35], where the selected adversary follows a predefined trajectory. This adversarial policy observes and acts in a closed-loop simulator alongside the ego policy. In this context, skill-based hierarchical policies are appealing approaches as they capture maneuvers at a higher abstraction, compared to the low-level actions of a simulator, corresponding more closely to how humans operate vehicles [26].
We build upon prior work [47], which utilizes expert demonstrations to extract paired observation and action sequences as state-conditioned “skills” which are then embedded using a Variational AutoEncoder (VAE). Additionally, a state-conditioned prior network is trained to map from a state to a useful location in the VAE’s latent space to be decoded into a reconstructed skill for the agent to follow.
@@ IV-C SEAL Implementation Details
For training and validating both the learned scoring function and skill spaces, we leverage the well-established Waymo Open Motion Dataset (WOMD) [8] dataset, as well as a subset of scenes therein labeled by Waymo as containing interacting agents. A further subset of 500 of these scenes has been used by prior work, and we henceforth refer to this set as WOMD-Normal [21, 48]. We split these scenes into 400 training and 100 evaluation scenarios.
@@
We implement $\pi_{\text{score}}$ as a VectorNet-style polyline encoder [51], followed by a multilayer perceptron decoder to the predicted values of $f_{\text{coll}}$ and $f_{\text{diff}}$. We use an MSE loss objective on the sum of the two values, ensuring equal weight to both predicted measures. For $\pi_{\text{adv}}$, we leverage the skill embedding framework from [47], with identical architectures and loss functions across our two parallel prior networks. We empirically set the hyperparameter $b$ in Equation 1 and Equation 2 to 8, use a skill time horizon of 10, and fix $K$ to 5 (consistent with CAT).
@@ V Experimental Setup
TABLE I: Ego performance on adversarially-perturbed (a, b, c) and unmodified, real-world (d, e) scenarios. WOMD-Normal are WOMD [8] scenes with basic interactive agents labeled by Waymo; WOMD-SafeShift-Hard refers to SafeShift-mined [16] real scenes in WOMD. Adversarially-perturbed scenes use WOMD-Normal as base scenarios, in both training and evaluation settings.
@@
Eval Scenario Type Ego Success ($\downarrow$) Realism WD ($\downarrow$) Yaw WD ($\downarrow$) Acc WD ($\downarrow$) Road WD ($\downarrow$) Crash WD ($\downarrow$)
WOMD-Normal, Real Scenes 0.60 0.04 0.12 0.02 0.03 0.00
WOMD-SafeShift-Hard, Real Scenes 0.41 0.05 0.12 0.04 0.04 0.00
WOMD-Normal, GOOSE-Gen [35] 0.43 0.33 0.12 0.60 0.48 0.12
WOMD-Normal, CAT-Gen [21] 0.30 0.14 0.12 0.31 0.07 0.07
WOMD-Normal, SEAL-Gen 0.31 0.09 0.12 0.16 0.05 0.05
@@ V-C Metrics
For evaluating generated scenario quality, we examine the induced ego Success rate, across all tested ego methods. We derive a realism metric based on distributional measures, similar to prior work  [12, 13, 32]. In particular, we utilize the Wasserstein distance (WD) over adversarial “profiles”—normalized histograms constructed from the adversary’s yaw rates, acceleration values, out of road rates, and rates of collisions with non-ego agents. All WD values are compared to profiles derived from the original $X_{\text{adv}}$ in $\mathbf{S}$. Furthermore, we perform a simple average over the four derived WD values to comprise an overall Realism meta-metric.
@@ VI Results
Downstream Performance. Our closed-loop training results are summarized in Table I. SEAL-trained policies average a $\mathbf{21.5\%}$ increase in Success rate relative to the top baseline in each evaluation setting, achieving a strong balance between Crash and Out of Road rates. While a baseline-trained policy may have slightly better performance on one failure type, it is achieved by sacrificing performance against the other.
Scenario Generation Quality. To directly assess scenario generation quality, we aggregate metrics in Table II, averaged over all ego methods. Note that WOMD-Normal and WOMD-SafeShift-Hard can have non-zero WD values due to early episode terminations from ego policies’ failures. Although CAT scenes induce a lower ego Success rate than SEAL scenes, SEAL scenes exhibit the highest Realism among scenario generation approaches, a $\mathbf{35.7\%}$ improvement, contributing to SEAL-trained policies’ superior downstream performance. We also showcase qualitative examples of the tested scenario generation approaches in Figure 3, against Replay. CAT and GOOSE perturbations both result in aggressive, colliding scenarios, while the SEAL perturbation creates a more nuanced near-miss scenario, reactively avoiding a direct collision.
Ablation Studies. To further investigate how different components of SEAL affect downstream training, we perform extensive ablation studies shown in Figure 2. We study the effect of our learned scoring function by comparing it to the heuristic, bounding box overlap approach used by CAT (Learned Obj and Heuristic Obj, respectively). Similarly, we compare our adversarial skill policy (Adv Skill Prior) with a benign prior variant (Benign Skill Prior) and a predefined trajectory following policy (TrajPred Adv). Note that the combination of TrajPred Adv with Heuristic Obj reduces to CAT. Our full SEAL approach performs the best across all evaluation settings; both the learned scoring function and adversarial skill policy appear to be needed.
@@ VII Conclusion
As autonomous driving (AD) systems advance, ensuring safety remains essential. While recent safety-critical scenario generation techniques show promise, they often lack the realism, reactivity, and nuance needed to provide strong training signals for closed-loop agents. We thus introduced Skill-Enabled Adversary Learning (SEAL) as a perturbation-based safety-critical scenario generation approach, combining a learned scoring function and an adversarial skill policy. In all test settings—across both real-world challenging scenarios and generated scenarios by SEAL and other SOTA methods—SEAL-trained policies achieved significantly higher success rates, with a more than 20% relative increase. Upon deeper analysis, SEAL-generated scenes contain less aggressive but more realistic adversaries, helping to explain the observed ego agent improvements. We argue that realism metrics, downstream task utility, and out-of-distribution evaluation settings are vital in assessing adversarially-perturbed scenarios.
While SEAL is quite effective, future improvements are still possible. Incorporating finer-grained objectives into the scoring function could enable more adaptive and controllable generation beyond safety criticality alone. Additionally, enhancing realism metrics to reflect human decision-making at the skill-level could provide deeper insights into scenario quality. We encourage future work to explore these topics.

### v2

@@ Abstract
Verification and validation of autonomous driving (AD) systems and components is of increasing importance, as such technology increases in real-world prevalence. Safety-critical scenario generation is a key approach to robustify AD policies through closed-loop training. However, existing approaches for scenario generation rely on simplistic objectives, resulting in overly-aggressive or non-reactive adversarial behaviors. To generate diverse adversarial yet realistic scenarios, we propose SEAL, a scenario perturbation approach which leverages learned objective functions and adversarial, human-like skills. SEAL-perturbed scenarios are more realistic than SOTA baselines, leading to improved ego task success across real-world, in-distribution, and out-of-distribution scenarios, of more than 20%. To facilitate future research, we release our code and tools: https://navars.xyz/seal/
@@ I Introduction
Scenarios are often curated in the form of large datasets of real-world recorded driving traces, providing a basis for assessing human behaviors and for training machine learning models [10, 11, 12]. AD subsystems are then asked to perform tasks such as forecasting the future motion of various road users or controlling the behavior of certain vehicles in a simulated reconstruction [13, 14, 15, 16]. However, the presence of critical scenarios in collected datasets is exceedingly low, a problem identified as the “curse-of-rarity” in autonomous driving [17, 18, 19]. Thus, programmatically generating safety-critical scenarios is necessary. To ensure that generated scenarios retain realistic properties, it is appealing to perturb the behavior of one or more agents in a principled way, rather than using first principles to painstakingly assemble a scenario from scratch [20, 21, 22, 1]. In this setting, one agent is referred to as the ego agent, while the modified background traffic participants are adversary agent(s), who attempt to attack the ego in some way.
@@
Therefore, in this paper, we propose and evaluate a method for Skill-Enabled Adversary Learning (SEAL), which yields significantly improved downstream ego behavior, in closed-loop training with safety-critical scenario generation. Our method addresses the identified limitations in prior art by introducing two novel components, as shown in LABEL:fig:overview. First, we introduce a learned objective function to anticipate how a reactive ego agent will respond to a candidate adversarial agent behavior. We quantify both collision closeness and induced ego behavior deviation, thus providing a broadened understanding of safety criticality. Second, we develop a skill-enabled, reactive adversary policy; in particular, inspired by human cognition, we leverage a hierarchical framework that is akin to how humans operate vehicles [27] and we create an adversarial prior that selects human-like skill primitives to increase criticality while maintaining realism.
Furthermore, we argue that safety-critical scenario generation should be evaluated based on behavior realism and usefulness for ego policy improvement, not just induced criticality. Prior work often assesses ego policies on generated scenarios where safety-critical behavior remains effectively in-distribution with respect to training data and heuristic perturbations [1, 28]. To address this, we build on recent scenario characterization work, SafeShift [18], to identify real (non-generated) but safety-relevant scenarios, enabling a more realistic, out-of-distribution evaluation. While in-distribution performance is informative, real-world performance on challenging scenes is ultimately most important.
@@
    We propose two novel techniques for safety-critical perturbation: (i) a learned objective function, to select candidate trajectories; and (ii) an adversarial skill-based, reactive policy for more realism in adversary behavior.
@@ II-A Scenario Generation in Autonomous Driving
Approaches for generating scenarios that reproduce the distribution of normal driving behavior have been extensively explored. Some methods ensure the diversity of generated traffic behavior [29, 30], while others aim for controllability through rule-based or language-driven specifications [31, 32, 33]. However, due to the rarity of safety-critical events in recorded data [17, 18, 19], other approaches have focused on directly generating corner-case scenarios by injecting adversarial behaviors. Earlier works in safety-critical scenario generation relied on gradient-based optimization approaches with access to vehicle dynamics [28, 25, 34], a limitation in model-free settings. Other methods, such as diffusion-based approaches [26, 35], are compute-intensive and impractical to be used in a closed-loop manner. Efficient methods like CAT [1] and GOOSE [2], which leverage trajectory prediction priors and reinforcement learning (RL) respectively, prioritize simple collision objectives and are non-reactive to the ego agent. In contrast, our approach efficiently generates reactive, nuanced adversarial behavior, providing a stronger training signal and improving closed-loop agent performance.
@@ II-B Robust Training and Evaluation in Autonomous Driving
Several techniques for robustifying AD policies against safety-critical and out-of-distribution scenarios have been explored. Formal methods, such as Hamilton-Jacobi (HJ) reachability, have been utilized in various driving tasks, but struggle with dimension scaling [36, 37]. Similarly, domain randomization has been used as a form of data augmentation (e.g., randomizing vehicle control parameters [38] or scenario initial states [39]) but requires excessive sampling to cover a sufficient domain size. Thus, adaptive stress testing [17, 40] and adversarial training have been increasingly used, either as a fine-tuning scheme [26, 25] or in a fully closed-loop training pipeline [1, 34], providing continuous feedback to an ego agent. However, these approaches still tend to optimize for naive collision objectives alone.
@@ IV Approach: Skill-Enabled Adversary Learning for Scenario Generation
To increase the criticality while maintaining the realism of a scenario, we propose the Skill-Enabled Adversary Learning (SEAL) approach for perturbation-based scenario generation. We share a similar insight to CAT [1] of utilizing a probabilistic trajectory prediction model, $\pi_{\text{gen}}$, to produce diverse candidate future paths that an adversary could take, conditioned on $\mathbf{S}$. However, there are several challenges with using this previously-explored approach of directly providing a future path for the adversary to follow from $\pi_{\text{gen}}(\mathbf{S})$:
@@
To address these challenges, we introduce two key components: a learned objective function to select trajectories more flexibly than heuristic methods, and an adversarial skill policy that enables more human-like and reactive behavior.
@@
IV-A Learned Objective Function
@@ IV-B Adversarial Skill Learning
We design a reactive policy $\pi_{\text{adv}}$, to guide the adversary’s behavior, unlike recent works [1, 2], where the selected adversary follows a predefined trajectory. This adversarial policy observes and acts in a closed-loop simulator alongside the ego policy. In this context, skill-based hierarchical policies are appealing approaches as they capture maneuvers at a higher abstraction, compared to the low-level actions of a simulator, corresponding more closely to how humans operate vehicles [27].
We build upon prior work, ReSkill [48], which utilizes expert demonstrations to extract paired observation and action sequences as state-conditioned “skills” which are then embedded using a Variational AutoEncoder (VAE). Additionally, a state-conditioned prior network is trained to map from a state to a useful location in the VAE’s latent space to be decoded into a reconstructed skill for the agent to follow.
@@ IV-C SEAL Implementation Details
For training and validating both the learned objective function and skill spaces, we leverage the well-established Waymo Open Motion Dataset (WOMD) [10] dataset, as well as a subset of scenes therein labeled by Waymo as containing interacting agents. A further subset of 500 of these scenes has been used by prior work, and we henceforth refer to this set as WOMD-Normal [1, 49]. We split these scenes into 400 training and 100 evaluation scenarios.
@@
We implement $\pi_{\text{score}}$ as a VectorNet-style polyline encoder [52], followed by a multilayer perceptron decoder to the predicted values of $f_{\text{coll}}$ and $f_{\text{diff}}$. We use an MSE loss objective on the sum of the two values, ensuring equal weight to both predicted measures. For $\pi_{\text{adv}}$, we leverage the skill embedding framework from [48], with identical architectures and loss functions across our two parallel prior networks. We empirically set the hyperparameter $b$ in Equation 1 and Equation 2 to 8, use a skill time horizon of 10 steps, and fix $K$ to 5 (consistent with CAT).
@@ V Experimental Setup
TABLE I: Ego performance on adversarially-perturbed (a, b, c) and unmodified, real-world (d, e) scenarios. WOMD-Normal are WOMD [10] scenes with basic interactive agents labeled by Waymo; WOMD-SafeShift-Hard refers to SafeShift-mined [18] real scenes in WOMD. Adversarially-perturbed scenes use WOMD-Normal as base scenarios, in both training and evaluation settings. Higher success rates and lower crash and out of road rates are better.
@@
Eval Scenario Type Ego Success ($\downarrow$) Realism WD ($\downarrow$) Yaw WD ($\downarrow$) Acc WD ($\downarrow$) Road WD ($\downarrow$)
WOMD-Normal, Real Scenes 0.60 0.06 0.12 0.02 0.03
WOMD-SafeShift-Hard, Real Scenes 0.41 0.07 0.12 0.04 0.04
WOMD-Normal, GOOSE-Gen [2] 0.43 0.40 0.12 0.60 0.48
WOMD-Normal, CAT-Gen [1] 0.30 0.17 0.12 0.31 0.07
WOMD-Normal, SEAL-Gen 0.31 0.11 0.12 0.16 0.05
@@ V-C Metrics
For evaluating generated scenario quality, we examine the induced ego Success rate, across all tested ego methods. We derive a realism metric based on distributional measures, following MixSim [14] and other prior work [15, 33]. In particular, we utilize the Wasserstein distance (WD) over adversarial “profiles”—normalized histograms constructed from the adversary’s yaw rates, acceleration values, and out of road rates. All WD values are computed via comparison to profiles derived from the original $X_{\text{adv}}$ in $\mathbf{S}$. Furthermore, we perform a simple average over the three derived WD values to compute an overall Realism meta-metric.
@@ VI Results
Downstream Performance. Our closed-loop training results are summarized in Table I. SEAL-trained policies average a $\mathbf{21.5\%}$ increase in Success rate relative to the top baseline in each setting, achieving a strong balance between Crash and Out of Road rates. While a baseline-trained policy may have slightly better performance on one failure type, it is achieved by sacrificing performance against the other.
Scenario Generation Quality. To directly assess scenario generation quality, we aggregate metrics in Table II, averaged over all ego methods. Note that WOMD-Normal and WOMD-SafeShift-Hard can have non-zero WD values due to early episode terminations from ego policies’ failures. Although CAT scenes induce a lower ego Success rate than SEAL scenes, SEAL scenes exhibit the highest Realism among scenario generation approaches, a $\mathbf{35.3\%}$ improvement, contributing to SEAL-trained policies’ superior downstream performance. We also showcase qualitative examples of the tested scenario generation approaches in Figure 3, against Replay. CAT and GOOSE perturbations both result in aggressive, colliding scenarios, while the SEAL perturbation creates a more nuanced near-miss scenario, reactively avoiding a direct collision.
Ablation Studies. To further investigate how different components of SEAL affect downstream training, we perform extensive ablation studies shown in Figure 2, as well as comparing against CAT as it is a slightly stronger baseline than GOOSE. We study the effect of our learned objective function by comparing it to the heuristic, bounding box overlap approach used by CAT (Learned Obj and Heuristic Obj, respectively). Similarly, we compare our adversarial skill policy (Adv Skill Prior) with a benign prior variant (Benign Skill Prior) and a predefined trajectory following policy (TrajPred Adv). Our full SEAL approach performs the best across all evaluation settings; both the learned objective function and adversarial skill policy appear to be needed.
@@ VII Conclusion
As autonomous driving (AD) systems advance, ensuring safety remains essential. While recent safety-critical scenario generation techniques show promise, they often lack the realism, reactivity, and nuance needed to provide strong training signals for closed-loop agents. We thus introduced Skill-Enabled Adversary Learning (SEAL) as a perturbation-based safety-critical scenario generation approach, combining a learned objective function and an adversarial skill policy. In all test settings—across both real-world challenging scenarios and generated scenarios by SEAL and other SOTA methods—SEAL-trained policies achieved significantly higher success rates, with a more than 20% relative increase. Upon deeper analysis, SEAL-generated scenes contain less aggressive but more realistic adversaries, helping to explain the observed ego agent improvements. We argue that realism metrics, downstream task utility, and out-of-distribution evaluation settings are vital in assessing adversarially-perturbed scenarios.
While SEAL is quite effective, further improvements are still possible. Incorporating finer-grained metrics into the objective function could enable more adaptive and controllable generation beyond safety criticality alone. Additionally, enhancing realism metrics to reflect human decision-making at the skill-level could provide deeper insights into scenario quality. We encourage future work to explore these topics.